OpenStudy (anonymous):

How to graph this function log with 3 base of x=x-2

OpenStudy (agreene):

\[\log_3x = x-2\] is what you wrote. that is not a function, I'm assuming there is a y somewhere in there

OpenStudy (anonymous):

Nope that is what my professor wrote it to me to solve it and it really confusing me because there is no Y...

OpenStudy (anonymous):

The equation that you wrote is exactly what I'm asking but it really gave me a hard time because there's no Y

OpenStudy (agreene):

well, then i suppose we should solve for x and see what happens, either things will cancel out or it will be a line at a given x,

OpenStudy (agreene):

I suppose if you wanted to solve this using analytical geometry, which might be the best way (only thing I can think to do without invoking some real creepy algebra) graph the left and the right separately, where they cross you will get the values for x that solves this (there are two) your two eqns would be \(y=\frac{\log x}{\log 3}\) and \(y= x-2\)

OpenStudy (agreene):

for reference: http://www.wolframalpha.com/input/?i=solve+log_3+x+%3D+x-2

OpenStudy (anonymous):

Thank you so much I will work on it...